[Brahms-l] Shift report 20040109 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Fri Jan 09 2004 - 08:19:45 EST
------------------------------------------------------------
Date: Fri Jan  9 08:19:44 2004, Shift: 08:00-16:00
Supervisor: Catalin Ristea, with: 
Log book#15 pages 27 - 27

Shift Summary:
------------------------------------------------------------
1. Beam from 1am that still continues.

2. Data taken: 

     runs 9487 - 350k 
          9488 - 475k 
          9489 -  50k 
          9493 - 140k 
     all data on trigger 6 with:
         - MRS@90 without TPM2 and D5 off.
         - FS@4 with D1=D2=D3=D4=1/sqrt(2)B.

3. No new problems with detectors.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jan 9 08:19:53 2004

This archive was generated by hypermail 2.1.8 : Fri Jan 09 2004 - 08:20:09 EST